OpenStudy (anonymous):

I have no idea where to even start on this problem.. Which of the following is the correct factored form of the given equation? 4x 2 - 25 = 0 a) (4x - 5)(x + 5) = 0 b) (2x - 5)(2x + 5) = 0 c) 2(x - 5)² = 0

OpenStudy (anonymous):

A^2 - B^2 = (A-B)(A+B)?

OpenStudy (anonymous):

what is 2^2 = ? 5^2 = ? Remember the difference of squares. \[a ^{2} - b ^{2} = (a+b)(a-b)\]

OpenStudy (anonymous):

4 and 25

OpenStudy (anonymous):

Good now if 4 is the square of 2 and 25 is the square of 5 can you write out the factors. (2x+5)(????) what is the other one?

OpenStudy (anonymous):

No... look at the general formula...(a+b)(a-b) I a = 2 and b = 5 write out the equation in factored form. I did one half you do the other.. (2x+5)(ax-b) ???what is ax-b?

OpenStudy (anonymous):

4x^2 - 25 = 0 (2x)^2 - (5)^2 = 0 now using a^2 - b^2 = (a+b)(a-b) (2x +5)(2x-5)=0

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

u r welcome...☺